If you can cut the lawn in 2 hours, and your spouse can do it in 3 hours, how long would it take if you work together? What assumptions (if any) did you make to arrive at your solution?
: If you can cut the lawn in 2 hours, and your spouse can do it in 3 hours, how long would it take if you work together? What assumptions (if any) did you make to arrive at your solution?

Answer by stanbon(18984)

If you can cut the lawn in 2 hours, and your spouse can do it in 3 hours, how long would it take if you work together?
---------------
You DATA:
time = 2 hr/job ; rate = 1/2 job/hr
----------------------
Spouse DATA:
time = 3 hr/job ; rate = 1/3 job/hr
-----------------------
Together DATA:
time = x hr/job ; rate = 1/x job/hr
-----------------------
EQUATION:
rate + rate = together rate
1/2 + 1/3 = 1/x
5/6 = 1/x
x = 6/5 = 1 1/5 hr = 1 hr. 12 minutes
========================================
Cheers,
Stan H.